# SDFs from Unoriented Point Clouds using Neural Variational Heat Distances

### v1

@@ 4.4 Application: Geometric Queries
We demonstrate that our approach is accurate enough to compute the union or intersection of the sublevel sets $[\phi<0]$ for given SDFs. For two sublevel sets $[\phi_{1}<0]$ and $[\phi_{2}<0]$ we have
@@
Figure 14 shows an example result. Note that in general, neither $\min\{\phi_{1},\phi_{2}\}$ nor $\max\{\phi_{1},\phi_{2}\}$ is an SDF [21].
@@ 5 Limitations
Our method has a few limitations. To ensure that we solve only first-order variational problems that are convex with respect to the function gradients, we train two separate networks, one for the heat diffusion time step $u^{\tau}$, and one for the SDF $\phi$. We also observe a moderate loss of detail, which is a common drawback of a neural representations. Additionally, singularities such as crease lines lead to locally larger errors in the SDF. The proposed method for computing the sets $\mathcal{B}^{\pm}$ does not generalize to shapes with interior voids or disconnected internal structures, albeit such cases are relatively uncommon.
@@ Appendix A Proofs
Here we give the proofs of Proposition 2.1 and Theorem 2.2.
Proposition 2.1.
By the trace theorem $u\mapsto\fint_{S}u\,\mathrm{d}a$ is a bounded linear functional on $H^{1}(\Omega)$. The quadratic form $u\mapsto\int_{\Omega}u^{2}+\tau|\nabla u|^{2}\,\mathrm{d}x$ is $H^{1}(\Omega)$-coercive and bounded. Thus, by the Lax-Milgram theorem, there is a unique weak solution $u^{\tau}$ of the Euler-Lagrange equation $0=\partial_{u}{\mathcal{E}_{\mathrm{MM}}}(u)$ (5) and this solution is the minimizer of the energy ${\mathcal{E}_{\mathrm{MM}}}$ (4). ∎
The following lemma will be used in the proof of Theorem 2.2.
@@
$\leq|A(\phi_{j},n_{k})-A(\phi,n_{k})|$
$\quad+|A(\phi_{j},n_{k}-n)|+|A(\phi,n_{k}-n)|$
$\leq|A(\phi_{j},n_{k})-A(\phi,n_{k})|$
$\quad+(\|\nabla\phi_{j}\|_{L^{2}}+\|\nabla\phi\|_{L^{2}})\|n_{k}-n\|_{L^{2}}$
@@
The first term on the right-hand side is weakly lower semicontinuous on $H^{1}(\Omega)$. By Lemma A.1 the functional $\phi\mapsto A(\phi,n^{\tau})$ is weakly continuous on $H^{1}(\Omega)$. Therefore, ${\mathcal{E}_{n}}$ is weakly lower semicontinuous.
@@ Appendix B Additional Experimental Details
Here, we provide some additional qualitative and quantitative evaluation and details on the used data. In table 3 we provide quantitative evaluation for the noisy and sparse input point cloud of the capped torus shape from fig. 11.
In fig. 15, we show Marching Cubes results for the zero-level sets of our neural SDF on the set of shapes used in the quantitative evaluation in fig. 9.
In fig. 16, we show the zero-level set results for some additional shapes. They were all computed with the same parameter set that is mentioned in the main text. Notably, no parameter tuning is required to reconstruct thin structures like the birds claws.
Finally, table 4 gives an overview over all considered models and the respective point cloud sizes. The point cloud for the capped torus experiment was created by ourselves.
Figure 15: Marching Cubes result for zero-level sets of our neural SDF. Models used in the quantitative evaluation in fig. 9.

### v2

@@ 4.4 Application: Geometric Queries
We demonstrate that our approach is accurate enough to compute the union or intersection of the sublevel sets $\phi<0]$ for given SDFs. For two sublevel sets $\phi_{1}<0]$ and $\phi_{2}<0]$ we have
@@
Figure˜15 shows an example result. Note that in general, neither $\min\{\phi_{1},\phi_{2}\}$ nor $\max\{\phi_{1},\phi_{2}\}$ is an SDF [22].
@@ 5 Limitations
Our method has a few limitations. To ensure that we solve only first-order variational problems that are convex with respect to the function gradients, we train two separate networks, one for the heat diffusion time step $u^{\tau}$, and one for the SDF $\phi$. We also observe a moderate loss of detail, which is a common drawback of a neural representations. Additionally, singularities such as crease lines lead to locally larger errors in the SDF. The proposed method for computing the sets $\mathcal{B}^{\pm}$ does not generalize to shapes with interior voids or disconnected internal structures, albeit such cases are relatively uncommon. For these challenging cases, our method can be extended by using Dipole Normal Propagation (DNP) [24] for point cloud orientation and then generalized winding numbers [18] to determine $\mathcal{B}^{\pm}$. In Figure˜16, we show a failure case of our box algorithm, while the extended pipeline succeeds—at the cost of increased computation time.
@@ 6 Conclusion and Future Work
Figure 16: Input point cloud with three holes of varying sizes rendered as a mesh (left) and the extracted zero-level sets using our method with the simple box algorithm with standard grid size (2nd from left), using a four times larger grid size (3rd from left), and using DNP [24] and generalized winding numbers [18] to compute the inside/outside separation. The standard grid size results in an unsigned distance function in the corpus region, as it is smaller than the large hole. The larger grid size causes incorrect orientation in small regions, such as the nose, due to a lack of inside points.
@@ Appendix A Proofs
Here we give the proofs of Proposition˜2.1 and Theorem˜2.2.
Proposition˜2.1.
By the trace theorem $u\mapsto\fint_{S}u\,\mathrm{d}a$ is a bounded linear functional on $H^{1}(\Omega)$. The quadratic form $u\mapsto\int_{\Omega}u^{2}+\tau|\nabla u|^{2}\,\mathrm{d}x$ is $H^{1}(\Omega)$-coercive and bounded. Thus, by the Lax-Milgram theorem, there is a unique weak solution $u^{\tau}$ of the Euler-Lagrange equation $0=\partial_{u}{\mathcal{E}^{\mathrm{mod}}_{\mathrm{MM}}}(u)$ (5) and this solution is the minimizer of the energy $\mathcal{E}_{\mathrm{MM}}$} (4). ∎
The following lemma will be used in the proof of Theorem˜2.2.
@@
$\leq|A(\phi_{j},n_{k})-A(\phi,n_{k})|+|A(\phi_{j},n_{k}-n)|+|A(\phi,n_{k}-n)|$
$\leq|A(\phi_{j},n_{k})-A(\phi,n_{k})|+(\|\nabla\phi_{j}\|_{L^{2}}+\|\nabla\phi\|_{L^{2}})\|n_{k}-n\|_{L^{2}}$
@@
The first term on the right-hand side is weakly lower semicontinuous on $H^{1}(\Omega)$. By Lemma˜A.1 the functional $\phi\mapsto A(\phi,n^{\tau})$ is weakly continuous on $H^{1}(\Omega)$. Therefore, $\mathcal{E}_{n}$} is weakly lower semicontinuous.
@@ Appendix B Additional Experimental Details
Figure 17: Marching Cubes result for zero-level sets of our neural SDF. Models used in the quantitative evaluation in Figure˜11.
@@
Here, we provide some additional qualitative and quantitative evaluation and details on the used data. In Table˜3 we provide quantitative evaluation for the noisy and sparse input point cloud of the capped torus shape from Figure˜13.
In Figure˜17, we show Marching Cubes results for the zero-level sets of our neural SDF on the set of shapes used in the quantitative evaluation in Figure˜11.
In Figure˜18, we show the zero-level set results for some additional shapes. They were all computed with the same parameter set that is mentioned in the main text. Notably, no parameter tuning is required to reconstruct thin structures like the birds claws.
Finally, Table˜4 gives an overview over all considered models and the respective point cloud sizes. The point cloud for the capped torus experiment was created by ourselves.